URGENT!!!!!!!!!!!!!!!!!!

A standard American 100-Watt lightbulb is rated for the power it dissipates when connected to a US-style 110-V wall outlet. What is the resistance of such a lightbulb measured in ohms?
Answer #1


Power P = V2 / R
so, R = V2 / P
Resistance in ohms = Voltage 2 / Power = 1102 / 100 = 12100 / 100 = 121 ohms